4.0g of neon gas (Ne) at an initial temperature of 300K interacts thermally with 12.0g of oxygen gas (O2) at an initial temperature of 600K. The molar mass of atomic neon is 20g/mol, the molar mass of atomic oxygen is 16g/mol.
A) What is the initial thermal energy of each gas?
B) What is the final thermal energy of each gas?
C) How much heat energy is transferred and in which direction?
D) What is the final temperature?
Answer #1

Given

Mass of Neon gas                                            mN = 4.0 g

Mass of Oxygen gas                                        mO= 12.0 g

Initial temperature of Neon gas                                TN = 300 K

Initial Temperature of Oxygen gas           TO= 600 K

Molar mass of Neon                                       M­N = 20 g/mol

Molar mass of Oxygen                                   MO = 16 g/mol

Universal gas constant                                   R = 8.31J/molK

Solution

Number of moles in 4 grams of Neon gas

nN = mN / M­N

nN = 4.0/20

nN = 0.2 mol

Number of moles in 12 grams of Oxygen gas

nO = mN / M­N

nO = 12.0/20

nO = 0.6 mol

i)

Initial energy of Neon

UN = 3nNRTN/2

UN = 3 x 0.2 x 8.31 x 300 / 2

UN = 747.9 J

Initial energy of Oxygen

UO = 3nORTO/2

UO = 3 x 0.6 x 8.31 x 600 / 2

UO = 4487.4 J

ii)

Using the conservation of energy principle

Final energy of Neon + Final energy of Oxygen = Initial energy of Neon + Initial energy of Oxygen

UN’+UO’ =UN + UO                             …….(1)                 

The energy exchange will stop when both the gases are at the same temperature. Let T’ be the final temperature

UN’ = 3nNRT’/2

UO’ = 3nORT’/2

UN’/ UO’ = nN/nO

UN’ = nN UO’ /nO

Substituting these values in equation (1) we get

(nN UO’ /nO) +UO’ =UN + UO

UO’ {( nN /nO)+1} = 747.9 + 4487.4

UO’ {( 0.2/0.6)+1} = 5235.3

UO’ (0.8/0.6) = 5235.3

UO’ = 5235.3 x 0.6 / 0.8

UO’ = 3926.5 J is the final energy of Oxygen

UN’ = nN UO’ /nO

UN’ = 0.2 x 3926.5 /0.6

UN’ = 1308.8 J is the final energy of Neon

iii)

Change in neon gas energy

ΔUN = UN’ - UN

ΔUN = 1308.8 - 747.9

ΔUN = 560.9 J

Since the value of ΔUN is positive the neon gas gained 560.9 J.

Hence 560.9 J energy is transferred from oxygen to neon

iv)

UN’ = 3nNRT’/2

1308.8 = 3 x 0.2 x 8.31 x T’ /2

T’ = 262.5 K is the final temperature